OpenStudy (anonymous):

Solve the proportion x over x - 4 = 4 over x - 4. What is the value of x?

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow {x \over x-4} = {4 \over x-4} }\) Correct?

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow x(x - 4) = 4(x-4) }\) We cross-multiply when we have to solve a proportion.

OpenStudy (anonymous):

Yes that's correct. How do you solve for x?

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow x^2 - 4x = 4x - 16 }\) \(\Large \color{purple}{\rightarrow x^2 =8x - 16}\) \(\Large \color{purple}{\rightarrow x^2 - 8x = -16 }\) \(\Large \color{purple}{\rightarrow x^2 - 8x + 16 = 0 }\) Do you know how to solve a quadratic equation?

OpenStudy (anonymous):

This is a proportion. I'l give you my choices. 1. x = 2 2. x = -2 and x = 1 3. x = -1 and x = 4 4. x = 4

Parth (parthkohli):

Okay, \(\Large \color{purple}{\rightarrow x^2 - 4x - 4x + 16 = 0 }\) \(\Large \color{purple}{\rightarrow x(x - 4) - 4(x - 4) }\) \(\Large \color{purple}{\rightarrow (x - 4)^2 = 0 }\) \(\Large \color{purple}{\rightarrow x - 4 = 0 }\) \(\Large \color{purple}{\rightarrow x = 4 }\) 4th option :D

OpenStudy (anonymous):

Thanks you so much! :) Do you think you could help me with one more?

Parth (parthkohli):

I can!

OpenStudy (anonymous):

Okay. :)

OpenStudy (anonymous):

A map has a scale of 2 km to 3 inches. Which ratio of measures forms a proportion with the map’s scale of km to inches?

OpenStudy (anonymous):

DO you want me to give you my choices?

Parth (parthkohli):

Yes, you have to! It's a proportion.

OpenStudy (anonymous):

1. 12 over 18 2. 12 over 24 3. 12 over 14 4. 34 over 36

Parth (parthkohli):

So, basically, the answer has to be \(\Large \color{purple}{\rightarrow {2x \over 3x} }\)

OpenStudy (anonymous):

12 over 18...?

Parth (parthkohli):

I find the first correct, because you take x as 6, and you get: \(\Large \color{purple}{\rightarrow {2*6 \over 3*6} = {12 \over 18} }\) YESSS!!

OpenStudy (anonymous):

Yay! :D I understand! :) Thanks so much!